OpenStudy (anonymous):

How would you solve 1/3 + 1/4 = 1/x? I got 12/7 but I'm not sure if that's the way it's supposed to be answered?

OpenStudy (anonymous):

It is correct.

Parth (parthkohli):

Multiply both sides by 12 which eliminates the denominators.

OpenStudy (anonymous):

12/7

OpenStudy (anonymous):

sorry for it

hartnn (hartnn):

\(\huge \frac{1}{3}+ \frac{1}{4}= \frac{4+3}{4*3}= \frac{7}{12}\) so x = 12/7

OpenStudy (anonymous):

I made 12x the common denominator which made the numerators 4x + 3x = 12 and then I added 4x + 3x to get 7x and divided 7 on both sides of the equation.

hartnn (hartnn):

that is also correct

OpenStudy (anonymous):

okay. thank you!